Find the real roots of f(x)=x^4-16
Answer by jim_thompson5910(28504)

f(x) = x^4-16


0 = x^4-16


x^4-16 = 0


(x^2-4)(x^2+4) = 0


(x-2)(x+2)(x^2+4) = 0


x-2=0, x+2=0, or x^2+4=0


x=2, x=-2, x=2i, or x=-2i


So the real roots are 2 and -2